OpenStudy (anonymous):

Find the derivative of f(x) = 2/x at x = -2.

OpenStudy (anonymous):

@Jhannybean @ivancsc1996

OpenStudy (ivancsc1996):

What formulas of derivation do you know?

OpenStudy (jhannybean):

\[\large f(x)=\frac{2}{x}\]\[\large f(x)=2x^{-1}\] for starters.

OpenStudy (ivancsc1996):

Do you know the formula to derive 1/x?

OpenStudy (jhannybean):

Now can you find the derivative?

OpenStudy (ivancsc1996):

Or the power formula?

OpenStudy (anonymous):

uhhhhhh Okay lemme see

OpenStudy (jhannybean):

to find the derivative here, you can follow the format \[\large\frac{nx^{n-1}}{(n-1)}\]

OpenStudy (anonymous):

-4^-1 = 1/4

OpenStudy (anonymous):

POWER RULE WITH RATIONAL EXPONENTS. DONE.

OpenStudy (anonymous):

Lol for confidence, genius, you get best response. Except I don't know what to do with that lol

OpenStudy (jhannybean):

\[\large x^{-1}\] your -1 is your n. \[\large \frac{-1\cdot x^{-1-1}}{-1-1}\]

OpenStudy (anonymous):

-1*x^-2 / 1

OpenStudy (jhannybean):

\[\large \frac{-x^{-2}}{-2}= \frac{1}{2x^2}\]

OpenStudy (anonymous):

oh -2!

OpenStudy (jhannybean):

Now you have a 2 multiplying this.

OpenStudy (anonymous):

-4^2 = 16 so then 1/16

OpenStudy (anonymous):

uhhh this is hard >:(

OpenStudy (jhannybean):

So you'll have \[\large 2\cdot \frac{1}{2x^2}= \frac{1}{x^2}\]

OpenStudy (jhannybean):

\[\large f'(-2) = \frac{1}{(-2)^2}=?\]

OpenStudy (anonymous):

2*1/16 = 1/8 = 1/4

OpenStudy (anonymous):

thats 1/4

OpenStudy (jhannybean):

good job!

OpenStudy (anonymous):

but jhanny somethings wrong. My options are only 1, 1/2, -1/2, and -1

OpenStudy (jhannybean):

O_o

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

idk what went wrong. I think i might have multiplied or something wrong

OpenStudy (jhannybean):

Oof.one minute. I'll writeout what i did.

OpenStudy (anonymous):

okay thx

OpenStudy (jhannybean):

\[f(x)=2x^{-1}\]\[f'(x) = 2(-1)x^{-2}= -2\cdot \frac{1}{x^2} =-\frac{2}{x^2}\]

OpenStudy (ivancsc1996):

Look at it this way. You need to find the rate of change of 2/x. This is:\[\frac{ d }{ dx }2/x\]Since 2 doesnt change, then you can take it out:\[\frac{ d }{ dx }2/x=2\frac{ d }{ dx }1/x\]Now, the formula for deifferentiating 1/x is \[\frac{ d }{ dx }1/x=-\frac{ 1 }{ x ^{2}} \rightarrow 2\frac{ d }{ dx }1/x=-\frac{ 2 }{ x ^{2} } \rightarrow ANS=-\frac{ 2 }{ (-2)^{2} }=-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

-2/4 = -1/2

OpenStudy (jhannybean):

Ah... wasnt supposed to get rid of the 2. my bad.

OpenStudy (primeralph):